Q1: Find Vc(t) for t> 0 using general solution. 2 Ω X + 10V 20V V.CO 0.15F 522
Answer #1

2 tlo 2n lov + 5n Velo 2ov 7 a. os= 2000 I (0) = 20-10 Torbart entot- otot 7 NOW, 20=2x10 +Vcio) => vc (0) = 20-20 3 Game . VFor t₃D: lov zu + + 5n Velos 2+5 I (0) = 10 = (19/7) Ame. VC(O) = 2 XI (Q) - 20 volt 7 So, for 1st order circuit we know, tyo